# Puzzle Similarity: A Perceptually-guided No-Reference Metric for Artifact Detection in 3D Scene Reconstructions

### v1

@@ Abstract
Modern reconstruction techniques can effectively model complex 3D scenes from sparse 2D views. However, automatically assessing the quality of novel views and identifying artifacts is challenging due to the lack of ground truth images and the limitations of no-reference image metrics in predicting detailed artifact maps. The absence of such quality metrics hinders accurate predictions of the quality of generated views and limits the adoption of post-processing techniques, such as inpainting, to enhance reconstruction quality. In this work, we propose a new no-reference metric, Puzzle Similarity, which is designed to localize artifacts in novel views. Our approach utilizes image patch statistics from the input views to establish a scene-specific distribution that is later used to identify poorly reconstructed regions in the novel views. We test and evaluate our method in the context of 3D reconstruction; to this end, we collected a novel dataset of human quality assessment in unseen reconstructed views. Through this dataset, we demonstrate that our method can not only successfully localize artifacts in novel views, correlating with human assessment, but do so without direct references. Surprisingly, our metric outperforms both no-reference metrics and popular full-reference image metrics. We can leverage our new metric to enhance applications like automatic image restoration, guided acquisition, or 3D reconstruction from sparse inputs.
@@ 1 Introduction
Image-based rendering and 3D reconstruction from a sparse set of 2D views has received ample attention in recent years, both for pure geometry reconstruction and radiance-field modeling. Classical approaches using simple triangulation and epipolar geometry through methods such as structure from motion (SfM) to produce sparse point-clouds of diffuse color [28]. Densifying these representations can be done explicitly [13]. Alternatively, one can learn continuous, implicit representations [22, 3, 25], normally modeled through some kind of multi-layer perceptron (MLP). A tangential problem to these efforts is the collection of 2D data, and the handling of corrupted, distorted, or simply incomplete sets of images from an object or scene we would like to model. Learning representations from very sparse inputs has been a widely studied topic as well [44, 4, 38, 5]; they normally leverage learned priors on large datasets, helping to fill in the gaps by enforcing 3D consistency and that resulting reconstructions follow natural image statistics. However, quantifying the quality of novel views from these reconstructions is still problematic. These views can contain artifacts due to the sparsity of the dataset their model was constructed from, and automatically identifying them helps with restoration (e.g. masking for image-based inpainters [32]) or simply to guide future data acquisition to fill the gaps [15]. Recent works have followed a bayesian approach to quantification of the uncertainty of an area belonging or not to a reconstructed model or scene [7], which could potentially be leveraged for simple artifact detection; however, they require implicit representations of the scene, with fundamental changes to the scene model, and are not practical for more general applications that require visual artifact identification outside of scene reconstruction, as well as incapable of detecting artifacts not arising from lack of coverage.
@@
To tackle this, we propose a novel approach for artifact detection that can be leveraged on any set of images without an encoded explicit or implicit model of the scene or object they depict. As opposed to visual difference predictors (VDPs) [18] (which require references) and no-reference quality metrics [23, 24] (which typically do not provide maps, but rather produce single values of overall quality) our approach provides visual artifact maps with no direct references. We leverage learned perceptual patch statistics from small clean datasets and compare them to the embdedded statistics of new images from a similar distribution (i.e. novel reconstructed views from a 3DGS [13] representation) to obtain artifact maps without direct references. We test our generated maps through a human experiment where we ask participants to manually identify artifacts and distortions in images to generate ground-truth data of visual artifacts. Our results show that our method agrees with human assessment, correlating better than both no-reference and full-reference metrics. To summarize, our contributions are the following:
@@
  • A novel no-reference visual quality/artifact identification metric particularly tailored for image-based rendering,
  • a novel dataset of human-assessed artifact and distortion identification to validate our metric,
@@
  • and an application on image restoration and 3D reconstruction enhancement that showcases the utility of our approach.
@@ 2 Related Work
3D reconstruction and Image-based rendering
@@
Reconstructing 3D objects or scenes from sparse sets of 2D observations is a fundamental problem in vision [18]. Particularly, in the context of novel view synthesis, the objective is to approximate the radiance field (i.e. 5D function encoding spatially varying radiance emission) of specific objects or scenes. Most methods however, differ either on the model used to encode the function, or the rendering procedure. Implicit approaches model the radiance field as a continuous function, approximated by a multi-layered perceptron (MLP) [22, 31]. Rendering is usually done via sampling the implicit volume using ray-marching [36], which provides spatially varying values of density and anisotropic color emission modeled through Spherical Harmonics. Improvements over this formula have tackled performance limitations, either by using more efficient sampling techniques [27, 8, 25] or by distilling the implicit space into explicit density and anisotropic appearance volumes [43, 39]. On the other hand, purely explicit models do not require any pre-training using implicit functions, and were originally Eulerian in nature [42]. Explicit models are easier to optimize, usually faster, and more interpretable, which can help in different tasks such as scene editing or animation. More recently, anisotropic Lagrangian approaches have found tremendous success [13]. However, these explicit methods have introduced some limitations of their own along the way. Methods like 3D Gaussian Splatting [13] can only model areas that are directly supervised, and degrade less gracefully than implicit counterparts when querying viewpoints substantially outside the training set coverage. Detecting artifacts arising from the lack of coverage is difficult due to the lack of reference images. Our method produces these masks via supervision on the training data solely, which can enable unsupervised restoration (automatic inpainting of the artifacts based on available context [6]) or simply automatically guide further image acquisition to complete the dataset efficiently [15].
Image Quality Metrics (IQMs)
@@
We can roughly categorize IQMs on their reliance on reference images, or the lack of it. Full-reference techniques traditionally input a reference image and its distorted counterpart and compute pixel or patch wise differences, which can either be pooled across an image to provide a single quality value [37, 49] or in some cases, kept as a pixel-wise map to visualize the location or source of the distortions. Classic examples also include mean-absolute error (MAE) or mean-squared error (MSE). Pixel-map metrics share similarities with visual difference predictors (VDPs), with the distinction usually being that VDPs leverage some explicit or implicit model of the human visual system (HVS) to align distortion predictions with human assessment. In some cases, it is also possible to transform usually pooled metrics like SSIM [37], FSIM [46] and LPIPS [49] into pixel maps, and we show how in Section 4.1.
No-reference visual quality metrics relinquish the necessity of direct references, and are most commonly learned, often relying on mean-opinion based supervised learning. Alternatively to leveraging datasets on human quality assessment [24, 12], they can also be supervised on extracted features from natural image statistics [23, 40], or even synthetic scores [41]. They are not usually capable of producing spatial quality maps, which makes it impossible to determine the source of reduced quality or localize the potential artifacts.
Some IQMs however, are capable of producing visual maps as well [26, 40, 12]; this makes them the closest to our work. The most relevant work in this space, PIQE [26] similarly attempts to quantify visual artifacts without relying on any supervision; rather, it extracts local features from image patches and quantifies quality as a measure of distortions in the patch, in accordance with a low-level human vision model. In contrast, our work does not make any assumptions on the HVS, rather leveraging the latent space of a model pretrained on natural images to measure the euclidean distance in feature space of candidate image patches to a limited set of image patches from a similar distribution (i.e., images from the same scene in the context of scene reconstruction). This enables us to obtain a higher level of alignment with human assessment.
@@
VDPs usually differ from IQMs in the more explicit integration of models of human vision, leveraging the naturally compressive nature of the human visual system [18] to predict perceived differences between images. Improvements over the original framework have increased its robustness by extending their applicability to high dynamic range imagery [20]; making them eccentricity and motion-aware [19]; and integrating perceived color [21]. VDP frameworks have seen use in foveated rendering applications [35] and perceptually-aware tone mappers [34]. In contrast to VDPs, our model relinquishes the necessity of direct references, while performing similarly to state of the art VDPs such as FovVideoVDP [19].
@@ 3 Our Method
Let us establish an analogy for our method: pretend each reference image is a puzzle with many puzzle pieces. To test if a new image is similar to our references, we would simply shuffle all pieces from all puzzles from our references and try to reassemble the test image only using those pieces. If the new image is very similar to the references, we should have enough puzzle pieces to compose the other image confidently. However, if the image holds regions very different from what we saw in the reference images, we would lack puzzle pieces to assemble this area, effectively leaving holes in the newly assembled puzzle (image). An overview of our approach through this analogy can be seen in Figure 1.
In our work, the puzzle pieces correspond to embedded image patches. In order to assess patch similarity, an obvious approach would involve computing the dot product between all patches; best-matching pieces would be recorded to create a similarity map. This simplistic approach, however, would hardly align with human assessment. Inspired by the close correlation between human quality judgment and latent CNN feature maps [33, 49], we employ a pre-trained convolutional network [16, 10, 30] to embed all the references, computing similarity in the latent feature space. Note that comparing feature map ”pixels” in a convolutional network is similar to comparing individual patches in the input domain; this is due to the locality of the sliding kernels when convolving. The patch size is dependent on the receptive field (showcased in Figure 2).
Figure 2: Receptive field of a multi-layer CNN. Note how one pixel in the last layer can be seen as an embedded patch of the input space.
@@
Choosing the right layers for embedding is essential to maximize the quality of the predicted spatial maps. While early layers feature small receptive fields and capture fine details, deeper layers have larger receptive fields and capture coarser features. This can be observed in Figure 3, where we showcase different VGG layers. It is essentially a trade-off between prediction granularity, accuracy and speed. We identified that combining multiple layers into our metric computation incorporates the various levels of abstraction and scales in a robust manner. We thus compute the weighted average of the three layers; we empirically found that halving the image resolution more than three times did not significantly improve our results as the scale becomes too small and the pool of reference vectors too little and specific to find good correspondences among novel images, even for well-reconstructed areas. This observation suggested that features from the layers before the third down-sampling were most useful for our cause.
(a) Rendering
(b) $2^{\text{nd}}$ VGG layer
(c) $3^{\text{rd}}$ VGG layer
(d) $4^{\text{th}}$ VGG layer
Figure 3: Puzzle Similarity computed only on a single VGG layer respectively. Note how the second layer has a much better resolution and generally colder colors, while the fourth layer is much smoother and features a wider range of values. Warm colors indicate the presence of artifacts or poor reconstruction quality.
@@
We empirically found that while VGG produces the most fine-grained maps, AlexNet and SqueezeNet still managed to perform similarly, with the key advantage of doing so at a substantially reduced computational cost. We opted for SqueezeNet as it aligned best with our test examples, specifically using layers $\ell\in\{2,3,4\}$ with the weights $w_{2}=0.67$, $w_{3}=0.2$, and $w_{4}=0.13$, which we found heuristically.
@@ 4 Results
We will now analyze how our method stacks against competing approaches for both full-reference and no-reference visual map prediction in the context of reconstruction and image-based rendering.
In order to quantify the correlation between all these maps and human assessment, we present a novel dataset on human artifact identification, which we manually collected and can be found here 111www.placeholder__\__fake__\__link__\__for__\__dataset.comto facilitate future research on the topic.
As for our method, for each different scene, we compute embeddings on their respective training dataset and leverage each of them respectively to compute visual quality maps on validation and artifact-ridden views, as explained in Section 3.
4.1 A Novel Artifact Identification Dataset
We created a dataset of human-perceived artifacts in 3D reconstructed views with corresponding ground truths collected through a user study. In order to generate images with typical reconstruction artifacts, we run 3D Gaussian Splatting [13] on twelve scenes from Mip-NeRF360 [3], Tank and Temples [14] and Deep Blending [9] datasets, using default parameters but withholding substantial amounts of training images. Omitting training views increases the chance of artifacts appearing in those withheld views, while still allowing us to have clean references. For each dataset, we selected three renderings that demonstrated a mix of well-reconstructed areas, strong artifacts, and subtle artifacts, resulting in 36 samples across 12 datasets.
Experiment details
We asked 23 participants to segment visible artifacts in each of the 36 sample images under controlled viewing conditions using the tool developed by Wolski et al. [1], which the authors kindly provided. We include details on the participants’ self-reported gender and age distributions in Supplementary material, as well as detailed viewing and display conditions. During the experiment, users had no undistorted, artifact-free references at their disposal and thus had to judge individual images at face value. They would then mark areas found to be unnatural or unappealing, creating a binary mask.
With the dataset, we can evaluate the agreement between human judgment and any metric output by simply averaging all binary masks to estimate the probability of each pixel being marked as an artifact. Figure 4 shows example renderings (a) alongside metric predictions (b)-(d) and their average human-produced mask (e).
We evaluate our method against both full-reference and no-reference IQMs, and the state-of-the-art VDP. To assess their alignment with human perception, we will correlate their maps with the human ratings from our dataset, which we described in Section 4.1.
4.2 Comparison with Full-Reference Metrics
Adapting pooled metrics to produce spatial quality maps
In terms of single-valued quality metrics, we selected commonly used reconstruction quality metrics (L1, L2, SSIM [37], LPIPS [49]). We adapt L1, L2 and SSIM by simply removing the pooling step. For LPIPS, after computing the distance in the embedding space, the metric already upsamples each feature map back to the original image resolution. We can then pool across maps but not along 2D image dimensions, allowing us to retain the map. We compare LPIPS with three base models: VGG, AlexNet, and SqueezeNet. We further compare with F LIP [2], a commonly found metric in quality assessment, particularly in the field of physically-based rendering; it already produces spatial quality maps by default. The no-reference metric CNNIQA [12] was applied on patches as described in their paper. We applied padding to avoid cropping the borders and upsampled the final map. PIQE [26] already produces three different kinds of maps that we averaged.
To correlate them to our human ratings, we first compute each metric map for each sample rendering and then compute the Pearson correlation coefficient (PCC) and Spearman’s rank correlation coefficient (SRCC). To account for the different domains of the compared metrics and possibly non-linear relations, we fit a 5-parameter logistic curve for a fair comparison as suggested by [1, 17, 47, 29]:
where $x$ is an individual quality score and $a_{1\dots 5}$ are the tunable parameters optimized using gradient ascent to maximize either PCC or SRCC.
@@ 4.2 Comparison with Full-Reference Metrics
Figure 4: Selection of image quality maps for artifact-ridden renderings from various scenes. The last column shows ground-truth human assessments from our collected dataset.

### v2

# Puzzle Similarity: A Perceptually-guided Cross-Reference Metric for Artifact Detection in 3D Scene Reconstructions

@@ Abstract
Modern reconstruction techniques can effectively model complex 3D scenes from sparse 2D views. However, automatically assessing the quality of novel views and identifying artifacts is challenging due to the lack of ground truth images and the limitations of No-Reference image metrics in predicting reliable artifact maps. The absence of such metrics hinders assessment of the quality of novel views and limits the adoption of post-processing techniques, such as inpainting, to enhance reconstruction quality. To tackle this, recent work has established a new category of metrics (Cross-Reference), predicting image quality solely by leveraging context from alternate viewpoint captures [53]. In this work, we propose a new Cross-Reference metric, Puzzle Similarity, which is designed to localize artifacts in novel views. Our approach utilizes image patch statistics from the input views to establish a scene-specific distribution, later used to identify poorly reconstructed regions in the novel views. Given the lack of good measures to evaluate Cross-Reference methods in the context of 3D reconstruction, we collected a novel human-labeled dataset of artifact and distortion maps in unseen reconstructed views. Through this dataset, we demonstrate that our method achieves state of the art localization of artifacts in novel views, correlating with human assessment, even without aligned references. We can leverage our new metric to enhance applications like automatic image restoration, guided acquisition, or 3D reconstruction from sparse inputs. Find the project page at https://nihermann.github.io/puzzlesim/.
@@ 1 Introduction
Image-based rendering and 3D reconstruction from a sparse set of 2D views has received ample attention in recent years, both for pure geometry reconstruction and radiance-field modeling. Classical approaches using simple triangulation and epipolar geometry through methods such as structure from motion (SfM) to produce sparse point-clouds of diffuse color [40]. Densifying these representations can be done explicitly [17]. Alternatively, one can learn continuous, implicit representations [28, 4, 32], normally modeled through some kind of multi-layer perceptron. A tangential problem to these efforts is the collection of 2D data, and the handling of corrupted, distorted, or simply incomplete sets of images from an object or scene we would like to model. Learning representations from very sparse inputs has been a widely studied topic as well [62, 5, 54, 6]; they normally leverage learned priors on large datasets, helping to fill in the gaps by enforcing 3D consistency and that resulting reconstructions follow natural image statistics. However, quantifying the quality of novel views from reconstructions is still problematic. These views can contain artifacts due to the sparsity of the training dataset, and automatically identifying them helps with restoration (e.g. masking for image-based inpainters [44]) or simply to guide future data acquisition to fill the gaps [19]. Recent works have followed a Bayesian approach to quantification of the uncertainty of an area belonging or not to a reconstructed scene [8], which could potentially be leveraged for simple artifact detection; however, they require implicit models, with fundamental changes to the scene model, and are not practical for more general applications that require visual artifact identification outside of scene reconstruction, as well as incapable of detecting artifacts not arising from lack of coverage.
@@
To tackle this, we propose a novel approach for artifact detection that can be leveraged on any set of images without an encoded explicit or implicit model of the scene or object they depict. As opposed to visual difference predictors (VDPs) [23] (which require references) and No-Reference quality metrics [29, 30] (which typically do not provide maps, but rather produce single values of overall quality) our approach provides visual artifact maps with no direct references. We leverage learned perceptual patch statistics from small clean datasets and compare them to the embedded statistics of new images from a similar distribution (i.e. novel reconstructed views from a 3DGS [17] representation) to obtain artifact maps without aligned references. We test our generated maps through a human experiment where we ask participants to manually identify artifacts and distortions in images to generate ground-truth data of visual artifacts. Our results show that our method agrees with human assessment, correlating better than No-Reference, Full-Reference, and state-of-the-art Cross-Reference metrics. To summarize, our contributions are the following:
@@
  • A novel Cross-Reference visual quality/artifact identification metric particularly tailored for 3D reconstruction,
  • a novel dataset of densely human-labeled artifact and distortion maps to fill the gap of validation benchmarks for Cross-Reference metrics,
@@
  • and applications on image restoration and 3D reconstruction enhancement that showcase our approach’s utility.
@@ 2 Related Work
2.1 3D reconstruction and Image-based rendering
@@
Reconstructing 3D objects or scenes from sparse sets of 2D observations is a fundamental problem in vision [23]. Particularly, in the context of novel view synthesis, the objective is to approximate the radiance field (i.e. 5D function encoding spatially varying radiance emission) of specific objects or scenes. Most methods however, differ either on the model used to encode the function, or the rendering procedure. Implicit approaches model the radiance field as a continuous function, approximated by a multi-layer perceptron [28, 43]. Rendering is usually done via sampling the implicit volume using ray-marching [48], which provides spatially varying values of density and anisotropic color emission modeled through Spherical Harmonics. Improvements over this formula have tackled performance limitations, either by using more efficient sampling techniques [34, 9, 32] or by distilling the implicit space into explicit density and anisotropic appearance volumes [61, 55]. On the other hand, purely explicit models do not require any pre-training using implicit functions, and were originally Eulerian in nature [60]. Explicit models are easier to optimize, usually faster, and more interpretable, which can help in different tasks such as scene editing or animation. More recently, anisotropic Lagrangian approaches have found tremendous success [17]. However, these explicit methods have introduced some limitations of their own along the way. Methods like 3D Gaussian Splatting [17] can only model areas that are directly supervised, and degrade less gracefully than implicit counterparts when querying viewpoints substantially outside the training set coverage. Detecting artifacts arising from the lack of coverage is difficult due to the lack of reference images. Our method produces these masks via supervision on the training data solely, which can enable unsupervised restoration (automatic inpainting of the artifacts based on available context [7]) or simply automatically guide further image acquisition to complete the dataset efficiently [19].
2.2 Image Quality Metrics (IQMs)
Full-Reference Metrics (FR)
@@ 2.2 Image Quality Metrics (IQMs)
FR techniques traditionally input a reference image and its distorted counterpart and compute pixel or patchwise differences, which can either be pooled across an image to provide a single quality value [52, 68] or kept as a pixel-wise map to visualize the location or source of the distortions. Classic examples also include mean-absolute error (MAE), mean-squared error (MSE), or peak signal-to-noise ratio (PSNR). In some cases, it is also possible to transform usually pooled metrics like SSIM [52], FSIM [64], MS-SSIM [51] and LPIPS [68] into pixel maps, and we show how in the Supplementary material. Metrics to evaluate renderings are more commonly yielding maps to localize poorly rendered areas[2].
@@
VDPs usually differ from IQMs in the more explicit integration of models of human vision, leveraging the naturally compressive nature of the human visual system [23, 24] to predict perceived differences between images. Improvements over the original framework have increased its robustness by extending their applicability to high dynamic range imagery [26]; making them eccentricity and motion-aware [25]; and integrating perceived color [27]. VDP frameworks have seen use in foveated rendering applications [47] and perceptually-aware tone mappers [46]. In contrast to VDPs, our model relinquishes the necessity of direct references, while performing similarly to state-of-the-art VDPs such as FovVideoVDP [25].
Multi-Modal-Reference Metrics (MMR)
The recent emergence of tasks like image captioning, text-to-image synthesis and text-based image editing demand multimodal quality assessment[39, 15, 3, 35]. Recent MMR metrics leverage joint embedding spaces, such as those learned by CLIP[38], and transformer-based architectures to evaluate the consistency between textual descriptions and visual content[11]. Others localize implausible or misaligned areas of Text-to-Image generated content[22] or assess general image quality[69, 50] from textual templates.
No-Reference Metrics (NR)
No-Reference visual quality metrics relinquish the necessity of direct references and are most commonly learned, often relying on mean-opinion-based supervised learning. Alternatively to leveraging datasets on human quality assessment [16, 14, 58], they can also be supervised on extracted features from natural image statistics [31, 29, 56, 30], or even synthetic scores [57]. Deep learning approaches like TRIQ [59] use Transformer architectures to regress Mean Opinion Scores (MOS) and MUSIQ [16] introduced multi-scale Transformers to alleviate resolution constraints of convolutional networks. Hybrid approaches use multi-modal architectures in conjunction with text templates to query image features like noisiness, sharpness, or contrast that can be translated to MOS[69, 50]. NR metrics that rely on global image features or operate in the frequency domain make it impossible to extract quality maps so we focus on variants that use patch- or pixel-wise statistics.
Localizing artifacts demands IQMs capable of producing visual maps. PIQE [33] attempts to quantify visual artifacts without relying on any supervision; rather, it extracts local features from image patches and quantifies quality as a measure of distortions in the patch, in accordance with a low-level human vision model. CNNIQA [14] was one of the first NR metrics to employ convolutional neural networks (CNN) to regress mean-opinion scores; more recently, PaQ-2-PiQ [58] uses region proposals to select quality-determining patches. PAL4VST [67] localizes specific artifacts that emerge from Image Synthesis Tasks through binary segmentation. In contrast, our work makes no assumptions on the HVS, rather leveraging the latent space of a model pre-trained on natural images to measure the cosine similarity in feature space of candidate image patches to a limited set of image patches from a similar distribution (i.e., images from the same scene in the context of scene reconstruction), rendering a higher level of alignment with human assessment.
Cross-Reference Metrics (CR)
CR methods utilize a set of unaligned reference images to assess the quality of a single query image belonging to the same distribution or ”scene”. They are applicable for novel view synthesis tasks where a scene is generally observed from multiple viewpoints yielding additional information. This category was recently established by Wang et al. [53] and their metric CrossScore serves as the main baseline for our evaluation. CrossScore relies on the cross-attention module[49] to correlate a query image with unaligned multiview images to predict quality maps. The maps predict the quality of 14x14 patches of the input image and are trained to mimic unpooled SSIM maps. However, SSIM has been repeatedly shown to be unaligned with human quality assessment and perception[36, 37, 68], which fundamentally limits the potential of CrossScore. To ensure that our method excels in localizing human apparent artifacts we collected a human-annotated dataset of novel view synthesis artifacts, and demonstrate state-of-the-art results on the task.
@@ 3 Our Method
Let us establish an analogy for our method: pretend each reference image is a puzzle with many puzzle pieces. To test if a new image is similar to our unaligned references, we would simply shuffle all pieces from all puzzles from our references and try to reassemble the test image only using those pieces. If the new image is very similar to the references, we should have enough puzzle pieces to compose the other image confidently. However, if the image holds regions very different from what we saw in the reference images, we would lack puzzle pieces to assemble this area, effectively leaving holes in the newly assembled puzzle (image). An overview of our approach through this analogy can be seen in Fig. 1.
In our work, the puzzle pieces correspond to embedded image patches. In order to assess patch similarity, an obvious approach would involve computing the dot product between all patches; best-matching pieces would be recorded to create a similarity map. This simplistic approach, however, would hardly align with human assessment if unprocessed patches were used. Inspired by the close correlation between human quality judgment and latent CNN feature maps [45, 68], we employ a pre-trained CNN [20, 12, 42] to embed all the references, computing similarity in the latent feature space. Note that comparing feature map ”pixels” in a CNN is similar to comparing individual patches in the input domain; this is due to the locality of the sliding kernels when convolving. The patch size is dependent on the receptive field (showcased in Fig. 2).
@@
Choosing the right layers for embedding is essential to maximize the quality of the predicted spatial maps. While early layers feature small receptive fields and capture fine details, deeper layers have larger receptive fields and capture coarser features. This can be observed in Fig. 3, where we showcase different VGG layers. It is essentially a trade-off between prediction granularity, accuracy and speed. We identified that combining multiple layers into our metric computation incorporates the various levels of abstraction and scales in a robust manner. We thus compute the weighted average of the three layers; we empirically found that halving the image resolution more than three times did not significantly improve our results as the scale becomes too small and the pool of reference vectors too little and specific to find good correspondences among novel images, even for well-reconstructed areas. This observation suggested that features from the layers before the third down-sampling were most useful for our cause.
@@
Figure 2: Receptive field of a multi-layer CNN. Note how one pixel in the last layer is an embedding of a patch of the input space.
(a) Rendering
(b) $2^{\text{nd}}$ VGG layer
(c) $3^{\text{rd}}$ VGG layer
(d) $4^{\text{th}}$ VGG layer
Figure 3: Puzzle Similarity computed on a single VGG layer. Note how the second layer has a finer resolution and mostly cold colors, while the fourth layer is much smoother and features a wider range of values. Warm colors indicate artifacts or poor reconstruction quality.
@@
We empirically found that while VGG produces the most fine-grained maps, AlexNet and SqueezeNet still managed to perform similarly; while doing so at a substantially reduced computational cost. We opted for SqueezeNet as it aligned best with our test examples, specifically using layers $\ell\in\{2,3,4\}$ with the weights $w_{2}=0.67$, $w_{3}=0.2$, and $w_{4}=0.13$, which we found heuristically.
@@ 4 Results
Figure 4: Selection of image quality maps for artifact-ridden renderings from various scenes. The last column shows ground-truth human assessments from our collected dataset. Note that our metric provides the finest resolution enabling better artifact segregation.
We will now analyze how our method stacks against competing approaches for both No-Reference and Cross-Reference visual map prediction in the context of reconstruction and image-based rendering. To quantify the correlation between all these maps and human assessment, we present a novel dataset on human artifact identification, which we manually collected and can be found here111www.placeholder__\__fake__\__link__\__for__\__dataset.com to facilitate future research on the topic. As for our method, for each different scene, we compute embeddings on their respective training dataset and leverage each of them respectively to compute visual quality maps on validation and artifact-ridden views, as explained in Sec. 3. CrossScore leverages the same training views for its quality map.
4.1 A Novel Artifact Identification Dataset
We created a dataset of human-perceived artifacts in 3D reconstructed views with corresponding ground truths collected through a user study. To generate images exhibiting typical reconstruction artifacts, we apply 3D Gaussian Splatting [17] to twelve scenes from the Mip-NeRF360 [4], Tanks and Temples [18], and Deep Blending [10] datasets. We use default parameters but intentionally withhold a significant portion of training images, which we later utilize for additional validation. By omitting these views during training, we increase the likelihood of artifacts appearing in the withheld images. For each dataset, we selected three renderings that demonstrated a mix of well-reconstructed areas, strong artifacts, and subtle artifacts, resulting in 36 samples across 12 datasets.
Experiment details
We asked 22 participants to segment visible artifacts in each of the 36 sample images under controlled viewing conditions using the tool developed by Wolski et al. [1], which the authors kindly provided. We include details on the participants’ self-reported gender and age distributions in the Supplementary, as well as detailed viewing and display conditions. During the experiment, users had no undistorted, artifact-free references at their disposal and thus had to judge individual images at face value. They would then mark areas found to be unnatural or unappealing, creating a binary mask. With the dataset, we can evaluate the agreement between human judgment and any metric output by simply averaging all binary masks to estimate the probability of each pixel being marked as an artifact. Fig. 4 shows example renderings (a) alongside metric predictions (b)-(d) and their average human-produced mask (e).
We evaluate our method against both No-Reference and Cross-Reference IQMs. To assess their alignment with human perception, we correlate their maps with the human ratings from our dataset, as described in Sec. 4.1. NR- and CR-IQMs are the only metrics that can detect artifacts without a direct reference, but the way we collected our dataset gives us access to a ground truth that is normally unavailable. This enables us to assess FR-IQMs and current VDPs too, although they are otherwise not suitable for the objective in question. We include extensive results in the Supplemental and show that, unlike any other competing metric, our metric even outperforms the best FR-IQMs on this benchmark, proving the general superiority of our method.
4.2 Evaluation
Table 1: Pearson and Spearman Correlation between NR-, and CR-IQMs and Human Perception per Dataset. The dashed line separates NR (above) from CR (below) metrics.
Pearson PAL4VST [67] 0.139 0.088 0.062 0.153 0.002 0.005 0.068 0.197 0.000 0.104 0.000 0.119
PaQ-2-PiQ [58] 0.194 0.420 0.432 0.138 0.305 0.428 0.613 0.452 0.534 0.667 0.220 0.424
PIQE [33] 0.266 0.267 0.255 -0.089 0.583 0.441 0.490 0.091 0.526 0.200 0.376 0.101
CNNIQA [14] 0.027 0.037 0.064 -0.068 -0.053 0.409 0.324 0.345 0.367 0.400 -0.133 0.005
[…]
PuzzleSim (ours) 0.594 0.565 0.618 0.461 0.609 0.675 0.768 0.636 0.505 0.642 0.717 0.593
Spearman PAL4VST [67] 0.083 0.080 0.027 0.111 0.004 -0.003 0.069 0.169 0.000 0.098 0.000 0.108
PaQ-2-PiQ [58] 0.214 0.495 0.435 0.009 0.261 0.176 0.616 0.329 0.476 0.696 0.152 0.329
PIQE [33] 0.209 0.409 0.291 -0.130 0.460 0.229 0.620 0.079 0.375 0.224 0.279 0.174
CNNIQA [14] -0.085 0.020 0.166 0.157 -0.130 0.255 0.375 0.380 0.316 0.395 -0.219 -0.066
[…]
PuzzleSim (ours) 0.468 0.393 0.382 0.499 0.428 0.428 0.658 0.601 0.307 0.540 0.548 0.440
To correlate metric outputs to our human ratings, we first compute each metric map for each sample rendering and then compute the Pearson correlation coefficient (PCC) and Spearman’s rank correlation coefficient (SRCC). To account for the different domains of the compared metrics and possibly non-linear relations, we fit a 5-parameter logistic curve for a fair comparison as suggested by [1, 21, 65, 41]:
where $x$ is an individual quality score and $a_{1\dots 5}$ are optimized through gradient ascent to maximize PCC or SRCC.